Problem solving involving inverse of sine 

 1.) Find the angle of inclination when balloon which is 170 m above the ground is connected to a meterological station by cable of length 200 m.

2) If a kite is 40ft. Off the ground and the string holding the kite is 42 ft long, what is the angle of elevation to the kite?


Problem involving cosine functions


1.) The top of a broken tree touches the ground at a distance of 15 feet from its base. If the length of the broken part is 17.4 feet, what angle does the broken end of the tree make with the ground?


2.) A 5 m ladder is leaned against a perpendicular wall such that its base is 2 m from the wall. Work out the angle between the ladder and the floor, giving your answer to two decimal places.


Problem involving inverse of tangent function


1.) Find the measures of angle of elevation of the top of the building which is 110 ft.high and Michael is 90 feel away from the base of the building.


2.) A ladder is placed against a 40 food high electric pole such that it touches the top of the pole. If the bottom of the ladder is 10 feet ayaw from the base of the pole, what angle does the bottom of the ladder make with the ground?


1
Expert's answer
2021-02-19T07:55:25-0500

solutions solving involving inverse of sine 

sin-1=170/200 = 58.21 degrees

sin-1= 40/42 =72.25 degrees


solutions involving cosine functions


cos-1 = 15/17 = 28.07 degrees

cos-1 =2/5 = 66.42 degrees


solutions involving inverse of tangent function


tan-1 = 110/90 = 50.71 degrees

tan-1 = 40/10 = 75.96 degrees